### v1

# Superpose Singular Features for Model Merging

Abstract
Model merging is a critical technique for combining the capabilities of multiple fine-tuned models without requiring additional training. While existing methods treat parameters as vectors, they overlook the intrinsic structure of linear transformation matrices - the core components that comprise the majority of model parameters. These matrices are fundamental to neural networks, mapping input representations to output features through linear combinations. Motivated by the linear representation hypothesis, we introduce task matrix and propose to Superpose Features from Task Matrix (SFTM), a novel approach that superposes features from individual task models into a merged model. SFTM employs singular value decomposition to identify feature bases of linear transformation matrices and solves a linear system to optimally combine them while preserving input-output mappings from individual task models. Extensive experiments on vision transformers and language models demonstrate that SFTM consistently outperforms existing methods, achieving superior performance and enhanced out-of-distribution generalization.
1 Introduction
Pre-trained models have become a cornerstone of modern deep learning, providing a powerful foundation for a wide range of tasks (Zhuang et al., 2020; Bommasani et al., 2021). To serve different tasks, fine-tuning the pre-trained models separately on different tasks is a common practice (Kenton and Toutanova, 2019; Hu et al., 2021). However, maintaining multiple fine-tuned models for different tasks can be cumbersome and resource-intensive, especially as the number of tasks and models continues to grow. Multitask learning (Sanh et al., 2021; Raffel et al., 2020b) is a promising approach to address this issue by training a single model on multiple tasks simultaneously, but it often requires a large amount of data and computational resources to achieve optimal performance (Pilehvar and Camacho-Collados, 2018; Phang et al., 2018; Fifty et al., 2021).
Recently, model merging (Matena and Raffel, 2022; Jin et al., 2022; Ilharco et al., 2022a; Yadav et al., 2024) offers a solution to this problem by consolidating multiple models into a single model that can perform well on multiple tasks effectively. This approach eliminates the need for maintaining multiple separate models, reduces storage and computational requirements, and potentially enables knowledge transfer between tasks. A foundational approach to model merging is parameter averaging (Choshen et al., 2022; Ilharco et al., 2022b; Wortsman et al., 2022), which simply takes the mean of corresponding parameters across models. This evolved into weighted averaging methods like Fisher Merging (Matena and Raffel, 2022) and RegMean (Jin et al., 2022) that assign different weights to parameters based on their importance. A key advancement came with the introduction of task vectors (Ilharco et al., 2022a) which represent the difference between fine-tuned and pre-trained weights and enable lightweight cross-task generalization. Various techniques (Huang et al., 2023; Yang et al., 2023; Yadav et al., 2024; Yu et al., 2024; Du et al., 2024) were subsequently developed to handle task conflicts by resetting redundant parameters, resolving sign conflicts, dropping and rescaling parameters, and managing parameter competition between tasks. However, these methods often rely on heuristic rules for merging parameters while overlooking the intrinsic structure of parameters.
Our method is motivated by the linear representation hypothesis (Mikolov et al., 2013; Arora et al., 2016; Olah et al., 2020). The linear representation hypothesis states that the representations of deep neural networks can be decomposed into combinations of feature vectors. This linearity enables neural networks to represent complex concepts through these combinations. Linear transformations (e.g., in linear and convolution layers) play a crucial role in this process, as they can activate specific features by measuring the similarity between a representation and a feature vector (via the inner product with the row vectors of the transformation matrix) and extract new features through linear combinations (via the weighted sum of the column vectors of the transformation matrix). These properties have been leveraged in recent years to interpret LLMs (Elhage et al., 2022; Bricken et al., 2023; Templeton et al., 2024) and design more efficient fine-tuning techniques (Hu et al., 2021; Zhang et al., 2023b). Consequently, in model merging, it is essential that the linear transformations in the merged model preserve the input-output mappings of those in the individual models for effective feature activation and extraction.
In this paper, we propose a method for merging linear transformation matrices while preserving the input-output mappings of individual models’ linear transformations. Our approach leverages Singular Value Decomposition (SVD) to merge matrices, treating left singular vectors as basis for output representations and right singular vectors as basis for input representations. The goal is to ensure the merged matrices preserve the directions of output features from original transformations when applied to corresponding input features across tasks. We show this problem reduces to solving a system of linear equations, enabling model merging in singular decomposition space while maintaining feature directions from various tasks. We validate our method on multiple benchmark datasets spanning vision and natural language processing tasks. The results demonstrate that our method consistently outperforms existing model merging techniques.
Notations
Throughout this paper, we use bold uppercase letters (e.g., $\mathbf{X}$) to denote matrices, bold lowercase letters (e.g., $\mathbf{x}$) to denote vectors, and regular lowercase letters (e.g., $x$) to denote scalars. For a matrix $\mathbf{X}$, its transpose is denoted as $\mathbf{X}^{\top}$. We denote task $i$ as $\mathsf{T}_{i}$ and its corresponding model parameters as $\theta_{i}$. For linear transformations, a linear transformation matrix in the model trained on task $i$ is represented as $\mathbf{P}_{i}$. The subscript ${}_{\text{pre}}$ indicates pre-trained parameters or matrices - for example, $\theta_{\text{pre}}$ and $\mathbf{P}_{\text{pre}}$ denote the parameters and linear transformation matrices of the pre-trained model, respectively.
2 Related Work
2.1 Model Merging of Fine-tuned Models
Model merging is a technique that combines multiple models into a single model to enhance performance or enable the model to perform multiple tasks. Previous studies have shown that averaging the weights of multiple models fine-tuned from the same pre-trained initialization is a promising approach for model merging. Fisher Merging (Matena and Raffel, 2022) advances beyond simple averaging by utilizing the Fisher information matrix to assess the importance of individual parameters, which are then weighted accordingly during the merging process. Similarly, RegMean (Jin et al., 2022) forms a linear regression problem with extra data for each layer and offers a closed-form solution for the merged model’s parameters by solving the regression problem.
Beyond parameter averaging, Task Arithmetic (Ilharco et al., 2022a) introduces task vectors and adding the task vectors of individual tasks to merge model, demonstrating their effectiveness and lightweight nature in facilitating cross-task generalization. Building on this concept, PEM Composition (Zhang et al., 2023a) extends the task arithmetic framework to merge LoRA (Hu et al., 2021), while Ties-Merging (Yadav et al., 2024) addresses task conflicts by resetting redundant parameters and resolving sign conflicts. These methods, however, use a single merging coefficient across all task vectors, which limits their flexibility. In contrast, Lorahub (Huang et al., 2023) and AdaMerging (Yang et al., 2023) use different coefficients for enhanced adaptability. Lorahub’s performance is limited as it only searches for coefficients at the task level, while AdaMerging requires complex training and unlabeled test datasets, making it applicable solely to classification problems. DARE (Yu et al., 2024) proposes drop and rescale as preprocessing steps when merging fine-tuned LLMs. PCB-Merging (Du et al., 2024) is a lightweight, training-free technique for model merging that balances parameter competition by intra-balancing parameter significance within tasks and inter-balancing parameter similarities across tasks, effectively enhancing performance across various scenarios and domains.
2.2 Linear Representation Hypothesis
The linear representation hypothesis states that neural networks encode information by summing up "feature vectors" (Mikolov et al., 2013; Arora et al., 2016; Olah et al., 2020), i.e., a layer of a network represents a set of features as a weighted sum of task-associated vectors. This hypothesis has been observed in various models, including word embeddings (Mikolov et al., 2013; Conneau et al., 2017), sentence embeddings (Bowman et al., 2015), Transformer language models (Meng et al., 2022; Hendel et al., 2023), and vision-language models (Trager et al., 2023; Perera et al., 2023). The hypothesis has been explioted in various fields, especially in probing (Alain and Bengio, 2018; Belinkov, 2022) and interpretability (nostalgebraist, 2020; Elhage et al., 2022; Bricken et al., 2023; Gao et al., 2024).
Especially, the linear representation hypothesis is prominently featured in recent works on mechanistic interpretablity of language models (Olsson et al., 2022; Elhage et al., 2022; Bricken et al., 2023; Templeton et al., 2024). In mechanistic interpretability, models are understood by decomposing them into interpretable components and understanding how these components interact. The linear representation hypothesis suggests that important features in neural networks are often represented through linear combinations of neuron activations, rather than complex nonlinear transformations. This hypothesis has gained significant empirical support through studies of language models. For instance, (Elhage et al., 2022) demonstrated a toy model that learned to resconstruct its input through linear combinations of its neurons. Furthermore, Bricken et al. (2023) showed that individual neurons in a language model can encode semantically meaningful features in a largely linear fashion, which is later sclaed to large language models (Templeton et al., 2024).
3 Method
3.1 Problem Formulation
We start with a set of tasks $\{\mathsf{T}_{1},\ldots,\mathsf{T}_{T}\}$ and various pre-trained models. The objective is to fine-tune these models either by updating all parameters or using parameter-efficient fine-tuning (PEFT) methods. The goal of model merging is to combine multiple fine-tuned models $\{\theta_{1},\ldots,\theta_{n}\}$ into a single model $\theta_{M}$ that can perform all tasks $\{\mathsf{T}_{1},\ldots,\mathsf{T}_{T}\}$ effectively without requiring access to training data. While existing methods treat model parameters as vectors to be merged, we focus on the crucial role of linear transformations, which comprise the majority of parameters and computations in neural networks. These transformations are essential as they project data into spaces where important features become more apparent, making information processing more efficient. To formally analyze these transformations during merging, we introduce the Task Matrix in Definition 3.1.
Definition 3.1 (Task Matrix).
The task matrix is defined by $\mathbf{M}_{i}=\mathbf{P}_{i}-\mathbf{P}_{\text{pre}}\in\mathbb{R}^{m\times n},$ where $\mathbf{P}_{i}$ is a linear transformation matrix for model of task $\mathsf{T}_{i}$, $\mathbf{P}_{\text{pre}}$ is the linear transformation matrix for the pre-trained model.
Remark 3.2.
Existing merging methods based on task vectors neglect the internal structure of linear transformation in neural networks, and merge parameters as vector, i.e., $\text{vec}(\mathbf{P}_{j})=\text{vec}(\mathbf{P}_{\text{pre}})+\gamma\text{vec}(\mathbf{M})$. These methods employ various heuristic rules to merge parameters, which requires many hyperparameter tuning.
To merge models, preserving the input-output mappings of linear transformations is essential for maintaining model capabilities. Our method focuses specifically on merging linear transformations in a way that retains their key input-output mappings through singular value decomposition, rather than treating all parameters equally as in previous approaches. Specifically, we focus on merging task matrices $\mathbf{M}_{i},i=1\ldots T$ into a single merged matrix $\mathbf{M}\in\mathbb{R}^{m\times n}$. For a input feature $\mathbf{r}\in\mathbb{R}^{n}$, we aim to ensure that the output of the merged model $\mathbf{M}\mathbf{r}$ maintains the same directional features as the outputs $\mathbf{M}_{i}\mathbf{r}$ from the original task-specific models, i.e., the merged model preserves the input-output mappings of the individual models. The final merged model is then constructed by adding the merged task matrix $\mathbf{M}$ to the pre-trained model with a scaling factor $\gamma$: $\mathbf{P}_{j}=\mathbf{P}_{\text{pre}}+\gamma\mathbf{M}$.
3.2 Merge Task Matrices to Superpose Singular Features
Identify Basis for Preserving Mappings
To keep the directional features of the task matrices $\mathbf{M}_{i}\mathbf{r}$ during merging, we first need to identify some basis features that is representative for the input and output representations. Singular value decomposition (SVD) of task matrices provides a natural way to obtain the basis features. Given the decomposition $\mathbf{M}_{i}=\sum_{k=1}^{r_{i}}\sigma_{i}^{(k)}\mathbf{u}_{i}^{(k)}\mathbf{v}_{i}^{(k)\top}$ where $\sigma_{i}^{(k)}$ is the $k$-th singular value, $\mathbf{u}_{i}^{(k)}\in\mathbb{R}^{m\times 1}$ and $\mathbf{v}_{i}^{(k)}\in\mathbb{R}^{n\times 1}$ are the $k$-th left and right singular vectors of $\mathbf{M}_{i}$, respectively. Therefore, the right singular vectors $\mathbf{v}_{i}^{(k)}$ form a basis for input representations, while the left singular vectors $\mathbf{u}_{i}^{(k)}$ form a basis for output representations. We named these basis features as singular features in our paper.
Except form the basis, singular features have several other properties that make them suitable for preserving input-output mappings during merging. First, the orthogonality of singular vectors within each task minimizes interference between features, enabling linear superposition of features in accordance with the linear representation hypothesis. Second, the decomposition reveals the model’s information processing mechanism - using inner products to detect input features and weighted sums to construct outputs. We leave the details of the good properties of singular features to the appendix.
Merge Task Matrices in Singular Space
Since we have the singular vectors as the basis of the input and output representations for linear transformation $\mathbf{M}_{i}$, the merged matrix $\mathbf{M}$ should be able to keep the directions of output features from original transformations when applied to corresponding input features across tasks. To accomplish this, we can merge the task matrices $\mathbf{M}_{i}$ in their singular spaces:
$\mathbf{M}=\sum\nolimits_{i=1}^{T}\sum\nolimits_{k=1}^{r_{i}}\alpha_{i}^{(k)}\sigma_{i}^{(k)}\mathbf{u}_{i}^{(k)}\mathbf{v}_{i}^{(k)\top},$ (1)
where $\alpha_{i}^{(k)}$ are the merging weights. To preserve feature directions during merging, we require that when the merged matrix $\mathbf{M}$ transforms an input representation, the output should maintain the same directional features as those produced by the individual task matrices $\mathbf{M}_{i}$. This leads to the following objective:
$\mathbf{u}_{i}^{(k)},\mathbf{M}\mathbf{v}_{i}^{(k)}\right>=\left<\mathbf{u}_{i}^{(k)},\mathbf{M}_{i}\mathbf{v}_{i}^{(k)}\right>,\quad\forall i,j,k.$ (2)
[…]
Solving Linear Systems to Merge Task Matrices
To solve (2), we rewrite it as a linear system for all tasks in $\{\mathsf{T}_{1},\ldots,\mathsf{T}_{T}\}$:
$\begin{split}&\mathbf{u}_{i}^{(k)\top}\!\!\left(\sum_{i^{\prime}=1}^{T}\sum_{k^{\prime}=1}^{r_{i}}\!\alpha_{i^{\prime}}^{(k^{\prime})}\sigma_{i^{\prime}}^{(k^{\prime})}\mathbf{u}_{i^{\prime}}^{(k^{\prime})}\mathbf{v}_{i^{\prime}}^{(k^{\prime})\top}\!\!\right)\!\mathbf{v}_{i}^{(k)}\!=\!\sigma_{i}^{(k)}\\ \Leftrightarrow&\sum_{i^{\prime}=1}^{T}\sum_{k^{\prime}=1}^{r_{i}}\!\!\sigma_{i^{\prime}}^{(k^{\prime})}\mathbf{u}_{i}^{(k)\top}\mathbf{u}_{i^{\prime}}^{(k^{\prime})}\mathbf{v}_{i^{\prime}}^{(k^{\prime})\top}\mathbf{v}_{i}^{(k)}\alpha_{i^{\prime}}^{(k^{\prime})}\!=\!\sigma_{i}^{(k)}.\end{split}$ (3)
For various tasks $\mathsf{T}_{i}$ and singular values, (3) forms a linear system with $r=\sum_{i=1}^{T}r_{i}$ variables $\alpha_{i^{\prime}}^{(k^{\prime})}$ and $r$ equations. The coefficient matrix of this linear system is the element-wise product of three matrices:
  • $\mathbf{\Sigma}=\mathbf{1}\otimes[\sigma_{i^{\prime}}^{(k^{\prime})}]_{i^{\prime},k^{\prime}}\in\mathbb{R}^{r\times r}$, where $\mathbf{1}\in\mathbb{R}^{r\times 1}$, $\otimes$ is the Kronecker product, and $[\sigma_{i^{\prime}}^{(k^{\prime})}]_{i^{\prime},k^{\prime}}$ is a row vector of singular values;
  • $\mathbf{U}=[\mathbf{u}_{i}^{(k)\top}\mathbf{u}_{i^{\prime}}^{(k^{\prime})}]_{i,k,i^{\prime},k^{\prime}}\in\mathbb{R}^{r\times r}$, where enumerating over $i,k$ forms the row index and enumerating over $i^{\prime},k^{\prime}$ forms the column index;
  • $\mathbf{V}=[\mathbf{v}_{i^{\prime}}^{(k^{\prime})\top}\mathbf{v}_{i}^{(k)}]_{i^{\prime},k^{\prime},i,k}\in\mathbb{R}^{r\times r}$, where enumerating over $i,k$ forms the row index and enumerating over $i^{\prime},k^{\prime}$ forms the column index.
Then, the coefficient matrix of the linear system is $\mathbf{\Sigma}\circ\mathbf{U}\circ\mathbf{V}$, where $\circ$ is the element-wise product. The linear equation (3) can be written as
$\mathbf{\Sigma}\circ\mathbf{U}\circ\mathbf{V}\boldsymbol{\alpha}=\boldsymbol{\sigma},$ (4)
where $\boldsymbol{\alpha}$ is the vector of variables $\alpha_{i^{\prime}}^{(k^{\prime})}$ and $\boldsymbol{\sigma}$ is the vector of singular values $\sigma_{i}^{(k)}$, with consistent indexing order over $i,k,i^{\prime},k^{\prime}$ across $\mathbf{\Sigma}$, $\mathbf{U}$ and $\mathbf{V}$. By solving the linear system in (4), we obtain the optimal weights $\boldsymbol{\alpha}^{*}$ for merging task matrices according to (1). The merged linear transformation matrix $\mathbf{P}$ is then obtained by adding the merged task matrix $\mathbf{M}$ to the pretrained matrix $\mathbf{P}_{\text{pre}}$ with a scaling factor $\gamma$: $\mathbf{P}=\mathbf{P}_{\text{pre}}+\gamma\mathbf{M}$. The algorithm for merging task matrices is summarized in Algorithm 1.
Algorithm 1 Merging task matrices: $\mathsf{{SFTM}}(\mathbf{M}_{i},i=1,\ldots,T)$
  Input: task matrices $\mathbf{M}_{i},i=1,\ldots,T$;
  Apply SVD to $\mathbf{M}_{i}$ to obtain $\sigma_{i}^{(k)}$, $\mathbf{u}_{i}^{(k)}$, and $\mathbf{v}_{i}^{(k)}$;
  Prepare $\mathbf{\Sigma}$, $\mathbf{U}$, and $\mathbf{V}$;
  Solve $\mathbf{\Sigma}\circ\mathbf{U}\circ\mathbf{V}\boldsymbol{\alpha}=\boldsymbol{\sigma}$ to obtain $\boldsymbol{\alpha}^{*}$;
  Obtain the merged task matrix $\mathbf{M}$ by (1);
  Return: merged task matrices $\mathbf{M}$.
3.3 Complete Algorithm
Except the linear transformation matrices, there are other parameters in the model that need to be merged, such as biases, normalization parameters, embeddings, and convolutional layers. For biases and embeddings, we merge them with task arithmetic, i.e., adding them task vectors together. For normalization parameters, we merge them by averaging the normalization parameters of the individual tasks because the normalization can be seen as a linear transformation with diagonal matrix. For convolutional layers, we convert them into linear transformations by reshaping the convolutional kernels into a matrix and then merge them with the same method as the task matrices.
We also follow Ties-Merging (Yadav et al., 2024) and apply a trimming step. This involves keeping only the top $\eta$ parameters by magnitude while setting the remaining parameters to zero. This preprocessing step helps reduce noise and focus on the most significant parameters during merging. We present the complete algorithm for merging models in Algorithm 2.
Algorithm 2 Complete Algorithm for Model Merging
1:  Input: finetuned models $\{\theta_{1},\ldots,\theta_{T}\}$, pretrained model $\theta_{\text{pre}}$
2:  Parameters: sparsity ratio $\eta$, scaling factor $\gamma$
3:  for each linear transformation layer do
4:     Extract task matrices $\mathbf{M}_{i}=\mathbf{P}_{i}-\mathbf{P}_{\text{pre}}$ for $i=1,\ldots,T$
5:     Keep top $\eta$ parameters by magnitude in each $\mathbf{M}_{i}$
6:     $\mathbf{M}\leftarrow\mathsf{{SFTM}}(\mathbf{M}_{i},i=1,\ldots,T)$Algorithm 1}
7:     Merged matrix $\mathbf{P}=\mathbf{P}_{\text{pre}}+\gamma\mathbf{M}$
8:  end for
9:  for each bias, embedding, normalization layer do
10:     Extract task vectors $\tau_{i}=\theta_{i}-\theta_{\text{pre}}$ for $i=1,\ldots,T$
11:     Keep top $\eta$ parameters by magnitude in each $\tau_{i}$
12:     if normalization layer then
13:        Merged parameters $\leftarrow\mathsf{mean}(\{\tau_{1},\ldots,\tau_{T}\})$
14:     else
15:        Merged parameters $\leftarrow\gamma\sum_{i=1}^{T}\tau_{i}$
16:     end if
17:  end for
18:  Return: merged model parameters
3.4 Discussion
Time Complexity
Because there are $T$ tasks matrices of size $m\times n$, SVD of these matrix takes the time complexity $O(Tmn^{2})$. To merge $T$ task matrices, the linear system has $r$ equations and variables to solve, which takes the time complexity $O(r^{3})$. Therefore, the overall time complexity of SFTM for merging $T$ task matrices is $O(Tmn^{2}+r^{3})$. See Section 4.3 for the time consumption of SFTM for various models.
Merging Task Matrices or Fine-tuned Matrices?
In this paper, our focus is on merging task matrices instead of fine-tuning matrices. We have discovered that fine-tuned linear transformation matrices tend to have more shared features across tasks. Therefore, merging the fine-tuned linear transformation matrix is not as efficient as task matrix because certain overlapping direction of singular vectors correspond to these common features. On the contrary, the singular vectors of the task matrix contain features that are more specific to individual tasks, making it a more effective approach for merging. See the results of merging fine-tuned matrices in Section 4.3.
4 Experiments
4.1 Experimental Setup
Evaluation Setup.
We evaluate SFTM across diverse fine-tuning methods, tasks, model architectures, and model size. For the fully fine-tuning, we evaluate vision tasks on ViT and NLP tasks on T5 following protocols of Ilharco et al. (2022a) and Yadav et al. (2024), respectively. For parameter-efficient fine-tuning (PEFT), we evaluate LoRA on GPT-2 for NLP tasks since it uses linear transformation matrices through low-rank adaptation, which aligns well with our approach, rather than vector parameters like IA3 (Liu et al., 2022) in previous research. To evaluate SFTM when model size scales up, we follow the setup of Du et al. (2024) to merge LLMs. Additionally, to evaluate the out-of-distribution robustness, we follow Yadav et al. (2024) to test it on NLP tasks.
Baseline Methods.
We compare SFTM against four established model merging approaches: (1) Averaging (Choshen et al., 2022; Wortsman et al., 2022), which computes the element-wise mean of individual models; (2) Task Arithmetic (Ilharco et al., 2022a), which merges by scaling and adding task vectors to the initial model; (3) Fisher Merging (Matena and Raffel, 2022), which approximates Fisher Information Matrix to weight parameters based on their importance for each task and combines the weighted parameters into the final merged model; (4) RegMean (Jin et al., 2022), which computes a closed-form solution to a least-squares regression problem that aims to minimize the distance between the merged model’s activations and the individual models’ activations. (5) Ties-Merging, which enhances merging by eliminating redundant parameters and resolving sign conflicts; and (6) PCB-Merging, which balances parameter competition through intra-task significance and inter-task similarity analysis. We also report results from individual fine-tuned models and a pre-trained model on all tasks.
4.2 Results
Table 1: Test set performance when merging ViT-B/32 and ViT-L/14 models on 8 vision tasks.
Task($\rightarrow$) Test Set Performance
Model Method($\downarrow$) Average SUN397 Cars RESISC45 EuroSAT SVHN GTSRB MNIST DTD
ViT-B/32 Pre-trained 46.3 61.7 54.7 58.5 51.2 29.1 27.4 45.6 42.1
Fine-tuned 90.5 75.3 77.7 96.1 99.7 97.5 98.7 99.7 79.4
Averaging 65.8 65.3 63.4 71.4 71.7 64.2 52.8 87.5 50.1
Fisher Merging 68.3 68.6 69.2 70.7 66.4 72.9 51.1 87.9 59.9
RegMean 71.8 65.3 63.5 75.6 78.6 78.1 67.4 93.7 52.0
Task Arithmetic 70.1 63.8 62.1 72.0 77.6 74.4 65.1 94.0 52.2
Ties-Merging 73.6 64.8 62.9 74.3 78.9 83.1 71.4 97.6 56.2
PCB-Merging 76.3 66.7 65.5 78.5 79.3 86.4 77.1 98.2 59.1
SFTM (ours) 77.4 65.4 62.0 81.4 90.4 84.1 77.1 95.6 62.9
ViT-L/14 Pre-trained 64.1 68.2 76.4 69.7 64.7 60.4 49.4 67.9 56.3
Fine-tuned 94.2 82.3 92.4 97.4 100 98.1 99.2 99.7 84.1
Averaging 79.6 72.1 81.6 82.6 91.9 78.2 70.7 97.1 62.8
Fisher Merging 82.2 69.2 88.6 87.5 93.5 80.6 74.8 93.3 70.0
RegMean 83.7 73.3 81.8 86.1 97.0 88.0 84.2 98.5 60.8
Task Arithmetic 84.5 74.1 82.1 86.7 93.8 87.9 86.8 98.9 65.6
Ties-Merging 86.0 76.5 85.0 89.4 95.9 90.3 83.3 99.0 68.8
PCB-Merging 87.5 76.8 86.2 89.4 96.5 88.3 91.0 98.6 73.6
SFTM (ours) 87.8 75.4 85.8 90.3 96.6 91.7 91.2 99.2 72.4
Fully Fine-tuned Vision Model: ViT Merging
We evaluate SFTM on vision tasks using two CLIP models (Radford et al., 2021) with ViT-B/32 and ViT-L/14 architectures (Dosovitskiy, 2020) as visual encoders. We use the released checkpoints from Ilharco et al. (2022a) that were fine-tuned on eight diverse classification tasks spanning remote sensing, traffic signs, and satellite imagery domains. These tasks include: Cars (Krause et al., 2013), DTD (Cimpoi et al., 2014), EuroSAT (Helber et al., 2019), GTSRB (Stallkamp et al., 2011), MNIST (LeCun, 1998), RESISC45 (Cheng et al., 2017), SUN397 (Xiao et al., 2016), and SVHN (Netzer et al., 2011). For all experiments, we keep the text encoder fixed and only merge the parameters of visual encoders.
As shown in Table 1, SFTM achieves state-of-the-art performance when merging fully fine-tuned ViT models, outperforming existing methods by 1.1% and 0.3% on average across 8 tasks for ViT-B/32 and ViT-L/14 architectures respectively. For ViT-B/32, SFTM shows particularly strong performance on RESISC45 and EuroSAT, with 2.9% and 11.1% improvements over the best baseline. For ViT-L/14, SFTM maintains high performance across all tasks, demonstrating effective preservation of task-specific features during merging.
Table 2: Merge fully-fine-tuned T5-base model with different methods
Test Set Performance
Method Average paws qasc quartz story_cloze wiki_qa winogrande wsc
Pre-trained 53.5 49.9 35.8 53.3 48.1 76.2 50.0 61.1
Fine-tune 81.0 91.4 95.5 79.1 79.6 95.2 62.8 63.6
Average 60.2 55.2 57.5 55.2 49.4 91.1 50.2 62.5
Fisher Merging 68.2 67.9 84.4 63.5 57.1 90.1 54.2 60.8
RegMean 71.5 76.2 92.8 62.6 63.6 89.4 57.4 58.3
Task Arithmetic 71.6 71.1 81.4 62.1 77.1 95.0 57.4 56.9
Ties-Merging 71.5 79.6 86.9 67.9 72.8 78.9 58.5 55.9
PCB-Merging 72.3 71.5 91.7 66.8 62.7 92.8 57.1 63.3
SFTM (ours) 73.8 78.2 89.4 61.7 74.9 94.6 55.8 61.7
Fully Fine-tuned NLP Model: T5 Merging
For NLP experiments, we evaluate on T5-base (Raffel et al., 2020a), an encoder-decoder transformer (Vaswani et al., 2017) pre-trained with masked language modeling. We finetune T5-base on seven diverse tasks spanning question answering (QASC (Khot et al., 2020), WikiQA (Yang et al., 2015), QuaRTz (Tafjord et al., 2019)), paraphrase identification (PAWS (Zhang et al., 2019)), sentence completion (Story Cloze (Sharma et al., 2018)), and coreference resolution (Winogrande (Sakaguchi et al., 2020), WSC (Levesque et al., 2012)). We use the code from Yadav et al. (2024) to finetune the model on these tasks. To reduce variance and ensure reliable evaluation, we report the experimental results of the average performance over different templates for each task.
As shown in Table 2, SFTM achieves state-of-the-art performance when merging T5-base models, outperforming existing methods by $1.5\%$ on average across 7 tasks. SFTM shows particularly strong performance on PAWS and Story Cloze, with improvements of $6.7\%$ and $12.2\%$ respectively over PCB-Merging, the previous best method. For Story Cloze, SFTM shows comparable performance to Task Arithmetic, while maintaining competitive performance on other tasks.
PEFT of NLP Model: LORA Adapters Merging
For PEFT experiments, we evaluate GPT-2 Medium model on LoRA adapters (Hu et al., 2021), which are task-specific adapters that are fine-tuned on NLP task datasets. These tasks include converting tables (E2E(Novikova et al., 2017)), knowledge graph (WebNLG(Gardent et al., 2017)) and structured data (DART(Nan et al., 2020)) to natural language. We use the released checkpoints from Hu et al. (2021) and evaluate the performance of the merged model on these datasets.
As shown in Table 3, SFTM outperforms existing baselines on all metrics. Specifically, SFTM show improvements of $3.2\%$ over TIES-merging for NIST of E2E, $4.8\%$ over PCB-Merging for CIDEr of DART.
Table 3: Merge LoRA Adapters of GPT-2 M with Different Methods. $\uparrow$ indicates higher is better, $\downarrow$ indicates lower is better.
Test Set Performance
E2E DART WebNLG
[…]
Pre-trained 0.2 0.58 1.5 5.2 0.002 0.2 2.0 152.4 0.15 2.0 179.1
LoRA 67.7 8.64 46.0 68.3 2.36 44.8 35.0 50.4 52.3 37.0 44.4
Average 63.4 8.00 40.8 66.6 2.01 40.0 32.0 53.7 43.9 32.0 49.2 5
Task Arithmetic 63.2 8.02 40.9 66.3 1.98 40.8 33.0 53.7 45.9 34.0 48.8 4
Ties-Merging 62.8 8.14 41.1 65.9 2.08 41.4 33.0 53.7 46.1 33.0 48.1 3
PCB-Merging 62.9 8.12 41.4 66.0 2.08 41.3 33.0 53.5 46.2 33.0 48.1 2
SFTM (ours) 64.1 8.40 42.2 66.5 2.18 41.6 33.0 54.1 47.1 34.0 48.0 1
Table 4: Results on the CMMLU, GSM8K, and Humaneval datasets.
Datasets
Method CMMLU GSM8K Humaneval Average
Chinese 38.6 2.3 13.4 18.1
Math 31.2 65.6 0 32.3
Code 33.3 0 17.1 16.8
Average 35.6 48.5 6.7 30.3
Task Arithmetic 35.4 46.1 9.8 30.4
Ties-Merging 36.5 53.4 12.8 34.3
PCB-Merging 36.4 52.3 16.5 35.1
SFTM (ours) 36.5 63.0 14.0 37.8
Large Model: LLM Merging
We evaluate model merging on three fine-tuned Llama-2-7B models (Touvron et al., 2023) focusing on different capabilities: Chinese language proficiency [^1] https://huggingface.co/LinkSoul/Chinese-Llama-2-7b, mathematical reasoning (Yu et al., 2023) [^2] https://huggingface.co/meta-math/MetaMath-7B-V1.0, and code generation (Rozière et al., 2023) [^3] https://huggingface.co/qualis2006/llama-2-7b-int4-python-code-18k. Each model’s performance was assessed using domain-specific benchmarks: CMMLU (Li et al., 2023) for Chinese language understanding, GSM8K (Cobbe et al., 2021) for mathematical reasoning, and HumanEval (Chen et al., 2021) for code generation abilities.
As shown in Table 4, SFTM achieves state-of-the-art performance across all three domains, improving overall performance by $2.7\%$ compared to the best baseline. The most significant improvement is observed in mathematical reasoning, where SFTM outperforms other methods by approximately $10\%$ on the GSM8K benchmark.
Table 5: Out-of-Distribution Generalization Performance of merged T5-base model
Out-of-Distribution Performance
Method Average cosmos_qa social_iqa quail wic copa h-swag
Pre-trained 31.1 21.9 18.8 24.1 65.6 43.8 12.5
Average 36.3 23.5 37.0 25.4 50.2 54.5 27.2
Task Arithmetic 37.0 21.9 36.8 25.5 49.5 61.4 26.6
Ties-Merging 37.1 22.2 37.8 24.9 51.6 62.2 26.5
PCB-Merging 37.3 23.0 38.1 24.6 52.2 59.0 27.2
SFTM (ours) 38.2 22.4 38.0 26.0 51.6 64.1 27.0
Out-of-Distribution Generalization
To evaluate out-of-distribution generalization, we test the merged T5-base model (previously trained on seven tasks) on six held-out tasks from the T0 mixture (Sanh et al., 2021): three question answering tasks (Cosmos QA (Huang et al., 2019), Social IQA (Sap et al., 2019), QuAIL (Rogers et al., 2020)), word sense disambiguation (WiC (Pilehvar and Camacho-Collados, 2018)), and two sentence completion tasks (COPA (Roemmele et al., 2011), H-SWAG (Zellers et al., 2019)). As shown in Table 5, SFTM achieves $0.9\%$ improvement over the best baseline on T5-base, demonstrating strong generalization capabilities to tasks outside the training distribution.
4.3 Additional Results and Analysis
(a) Task Vector Matrice vs Fine-tune Matrice
(b) Number of tasks
(c) Various $\eta$ at $\gamma$ = 0.6
(d) Various $\gamma$ at $\eta$ = 0.2
Figure 1: Performance of SFTM with different hyperparameters , number of tasks, and removing singular features.
Task Matrices versus Fine-tuned Matrices
We compare merging task matrices versus fine-tuned matrices directly on ViT-B/32, ViT-L/14 and T5-base models. As shown in Figure 1(a), merging task matrices consistently outperforms merging fine-tuned matrices across all architectures by a large margin. This validates our discussion in Section 3.4 that task matrices better isolate task-specific features compared to fine-tuned matrices, making them more effective for model merging.
Number of Tasks
We vary the number of tasks to evaluate the performance of SFTM. We perform experiments on the ViT-B/32. Following Du et al. (2024), We normalize the performance on each task to fine-tuned performance and then average the normalized performance across all chosen tasks. As shown in Figure 1(b), comparing to other methods, the performance of SFTM shows less decrease with the increasing number of tasks, which indicates that preserving the linear mapping of singular vectors are effective for model merging.
Hyperparameter Sensitivity
We analyze the sensitivity of SFTM to the hyperparameters $\eta$ and $\gamma$ on the ViT-B/32. Three model merging methods are compared in the experiments: SFTM, Ties-Merging, and PCB-Merging. We follow the experiment setup of previous work (Du et al., 2024) to analyze the effect of two parameters. The first parameter is the scaling factor $\gamma$ , which is used to scale the merged task vector. For SFTM and Ties-Merging, the second hyperparameter is the top-k percentage $\eta$, indicating that the $\eta\%$ parameters in task vector with the highest absolute value are retained. For PCB-Merging, the second hyperparameter is mask ratio $\eta$, indicating that the $\eta\%$ parameters in parameter competition balancing matrix with the biggest absolute value will be retained. We vary the hyperparameters $\gamma$ with step size 0.1 from 0.5 to 1.2, and $\eta$ with step size 0.05 from 0.1 to 0.5.
First, in Figures 1(c) and 1(d), SFTM achieves optimal performance with $\gamma<0.8$ and $\eta<0.4$ on ViT-B/32. Second, performance varies with both hyperparameters. For sparsity ratio $\eta$, performance first improves then declines as it increases. When $\eta$ is too low, important features are filtered out as most parameters are set to zero. When $\eta$ is too high, noise in the task matrices affects the quality of singular vectors. For scaling factor $\gamma$, we observe a similar trend - performance initially improves then deteriorates with increasing values. A small $\gamma$ diminishes the magnitude of singular vectors, while a large $\gamma$ may amplify interference between tasks.
5 Conclusion
In this paper, We present SFTM, a novel model merging approach that preserves task-specific features in linear transformations through singular value decomposition. Extensive experiments demonstrate that SFTM consistently outperforms existing methods across different architectures and tasks. The success of SFTM demonstrates the value of leveraging parameter structure in model merging, rather than treating all parameters as vectors. Our work establishes a foundation for future research on model merging that leverages the mathematical properties and structural roles of different parameter types.